F(x) = x2 + 9x - 4


1
Expert's answer
2022-03-09T08:16:21-0500

Solution


"f(x)=x^2+9x-4"


To solve this quadratic equation "x^2+9x-4=0", we use the quadratic formula


"x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\"


"a=1, b=9, c=-4"


"x = \frac{{ - \left( 9 \right) \pm \sqrt {{{\left( 9 \right)}^2} - 4\left( 1 \right)\left( { - 4} \right)} }}{{2\left( 1 \right)}}\"


"x = \frac{{ - 9 \pm \sqrt {81 + 16} }}{2}\"


"x = \frac{{ - 9 + \sqrt {97} }}{2},\,\,\,\,x = \frac{{ - 9 - \sqrt {97} }}{2}\"


"x=0.424, x=-9.424"


Hence the two solutions are


"x=0.424, x=-9.424"


The same has been shown in the following plot.